# Test-time regression: a unifying framework for designing sequence models with associative memory

### v1

@@ Abstract
resulting in an offset softmax attention as the output: $m_{(1,\mathbf{q}),\,t}(\mathbf{q})=\mathbf{M}^{(0)}=\bar{\mathbf{v}}-\mathbf{M}^{(1)}\bar{\mathbf{d}}$. This alternate form comes from the fact that if we already know $\mathbf{M}^{(1)}$, then we can solve the $p=0$ case with an alternate regression problem with preprocessed data $\{(\mathbf{k}_{i},\mathbf{v}_{i}-\mathbf{M}^{(1)}\mathbf{d}_{i})\}_{i=1}^{t}$, resulting in $\mathbf{M}^{(0)}$ being a normalized weighted sum of $\mathbf{v}_{i}-\mathbf{M}^{(1)}\mathbf{d}_{i}$. Unlike standard softmax attention (the $p=0$ case), which only considers the interaction between keys and queries, the $p=1$ case in Equation 37 also accounts for the covariance between keys through the inverse term, similar to how linear regression improves upon linear attention from before.
@@
Unfortunately, directly computing Equation 37 for $\mathbf{q}\in\{\mathbf{q}_{1},\ldots,\mathbf{q}_{T}\}$ in parallel to fully leverage hardware-acceleration is impractical for all but the smallest sequence problems. Materializing $\{\mathbf{X}_{t}\}_{t=1}^{T}$ requires $\mathcal{O}\left(T^{2}{D_{k}}\right)$ space, and materializing $\{\mathbf{X}_{t}{}^{\top}\mathbf{W}_{t}\mathbf{X}_{t}\}_{t=1}^{T}$ requires $\mathcal{O}\left(T{D_{k}}^{2}\right)$ space. Computing $\{\mathbf{X}{}^{\top}_{t}\mathbf{W}_{t}\allowbreak\mathbf{X}_{t}\}_{t=1}^{T}$ requires $\mathcal{O}\left(T^{2}{D_{k}}\right)$ time, and solving a batch of $T$ linear systems requires $\mathcal{O}\left(T{D_{k}}^{3}\right)$ time. These challenges makes the current form computationally impractical for all but the smallest sequence problems. Developing hardware-aware optimizations to compute Equation 37 similar to FlashAttention (Dao et al., 2022) is a promising future direction that would make higher order attention practical.
We compare our $p=1$ generalization of softmax attention against three other models on a non-stationary online regression problem. The keys are generated via a non-stationary (switching) autoregressive process:
@@
with ${D_{k}}=64$ and $T=256$. Each $\bm{\varepsilon}_{t}\sim\operatorname{Normal}\left(0,\bm{\Sigma}\right)$ where $\bm{\Sigma}$ has close to unit diagonals and non-zero off-diagonals. The choice of a non-identity $\Sigma$ creates instantaneous correlation between the dimensions of the keys so that $\mathbf{K}{}^{\top}\mathbf{K}\neq\mathbf{I}$. We let $\mathbf{v}_{t}=\lVert\mathbf{k}_{t+1}\rVert_{2}$, resulting in a next-token prediction task. Through the regression-memory correspondence, we directly apply each of these test-time regression layers to the key-value pairs without any learnable parameters; we simply pass the sequences through each layer via a single forward pass. We plot the one-step out-of-sample test loss $\lVert\mathbf{v}_{t+1}-{m}_{t}(\mathbf{k}_{t+1})\rVert_{2}^{2}$ in Figure 2. During the initial $T/4$ timesteps, the dynamical system has a faster decay time with more rapid changes, making it less predictable. In the latter $43T/$ 4 timesteps, the system slows down and is more predictable. However, unweighted recursive least squares is unable to adapt to this change since it cannot discount datapoints from before the transition; as a result, it maintains a high error throughout the latter timesteps. In contrast, the nonparametric regressors are able to adapt quickly to the transition since earlier keys are dissimilar to the later keys. Interestingly, increasing our nonparametric softmax regressor from $p=0$ to $p=1$ allows the model to better adapt to both the initial fast-changing period and the sudden regime change. As expected, linear attention performs the worst since $\mathbf{K}{}^{\top}\mathbf{K}\neq\mathbf{I}$ in our dynamical system [^5] We used a standard output normalization for linear attention (Qin et al., 2022) to prevent the output from scaling linearly with the sequence length..
@@
Section 4 Constructing effective key-value pairs
@@
Having shown that models derived via regression are able to perform associative recall, we now discuss the importance of constructing test-time key-value pairs that are pertinent to the task at hand. This echoes an unchanging principle in machine learning: a well-designed model is only as effective as the data it processes. Historically, query-key-value sequences were constructed via a linear projection on the corresponding timestep: $\mathbf{q}_{t}=\mathbf{W}_{Q}\mathbf{x}_{t},\,\mathbf{k}_{t}=\mathbf{W}_{K}\mathbf{x}_{t},\,\mathbf{v}_{t}=\mathbf{W}_{V}\mathbf{x}_{t}$ (Vaswani et al., 2017). Our framework naturally generalizes to the multihead case (Shazeer, 2019, Ainslie et al., 2023), corresponding to setting up parallel test-time regression problems, possibly with shared inputs or outputs.
Recent works on recurrent models, starting with Fu et al. (2022), have identified the importance of applying a “short convolution” to construct the query-key-value sequences, which can express the “shift” operation that induction heads learn (Olsson et al., 2022, von Oswald et al., 2023, Fu et al., 2022). The short convolution is a crucial component of purely recurrent language models, resulting in severe performance drops if removed (Yang et al., 2024b, c, a, Sun et al., 2024). In the case of transformers or model that combine recurrence with softmax attention (De et al., 2024), an explicit short convolution is not needed since earlier softmax attention layers can learn the behavior of a short convolution as empirically observed by von Oswald et al. (2023). Here we show that this short convolution is crucial to associative recall, creating bigram-like key value pairs.
One short conv is all you need for associative recall.
@@
Consider an abstracted version of multi-query associative recall (MQAR), a standard associative recall task introduced by Arora et al. (2023), in which a model must be able to recall multiple key-value pairings. In this task, there are $P$ pairs of unique and consistent cue-response [^6] We use the “cue” and “response” terminology to avoid confusion with “keys” and “values” in the context of test-time regression. tokens $\{(\mathbf{u}_{j},\mathbf{v}_{j})\}_{j\in[P]}$ with each cue $\mathbf{u}_{j}$ mapping one-to-one to its response $\mathbf{v}_{j}$. The model is given a contextual sequence of randomly drawn, possibly repeated, cue-response pairs followed by a previously seen cue: $(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T-1},\mathbf{x}_{T},\mathbf{x}_{T+1})=(\mathbf{u}_{i_{1}},\mathbf{v}_{i_{1}},\ldots,\mathbf{u}_{i_{T/2}},\mathbf{v}_{i_{T/2}},\mathbf{u}_{i_{j}})$. $\mathbf{u}_{i_{j}}$ is a cue that appeared earlier (the $j$th pair) and the model is expected to output the paired response $\mathbf{v}_{i_{j}}$ at timestep $T+1$.
We now show that a single test-time regression layer (e.g. as simple as linear attention) with one short convolution is sufficient to solve MQAR, without any parameters other than the embeddings, as long as it is provided with the appropriate key-value pairs to regress. To solve this task, it suffices to test-time memorize all the bigram pairing $\{(\mathbf{u}_{j},\mathbf{v}_{j})\}$. Given $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$, we causally construct our test-time regression dataset as follows. Let the keys be constructed via a short convolution $\mathbf{k}_{t}=w_{0}\mathbf{x}_{t}+w_{1}\mathbf{x}_{t-1}=\mathbf{x}_{t-1}$ where $w_{0}=0,w_{1}=1$. Let the queries and values be the same as the inputs: $\mathbf{q}_{t}=\mathbf{v}_{t}=\mathbf{x}_{t}$. At timestep $T+1$, we have $\mathbf{q}_{T+1}=\mathbf{x}_{T+1}=\mathbf{u}_{i_{j}}$. The output of a linear attention layer is then $\mathbf{y}_{T+1}=\sum_{t=1}^{T+1}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}\mathbf{q}_{T+1}=\sum_{t=1}^{T+1}\mathbf{x}_{t}\mathbf{x}_{t-1}{}^{\top}\mathbf{u}_{i_{j}}$ When the embedding space is large enough, we can set the embeddings such that all tokens are orthonormal. Then $\mathbf{x}_{t-1}{}^{\top}\mathbf{u}_{i_{j}}$ is nonzero only when $\mathbf{x}_{t-1}=\mathbf{u}_{i_{j}}$, simplifying the output to a sum of only tokens that are followed by $\mathbf{u}_{i_{j}}$. Since the cue-response map is one-to-one, the remaining tokens are all $\mathbf{v}_{i_{j}}$, producing the output $\mathbf{y}_{T+1}\propto\mathbf{v}_{i_{j}}$, solving the task.
@@
Typical evaluations of models on the MQAR task look at model performance with respect to the length of the sequence, possibly varying the model capacity. A few examples include Figure 2 of Arora et al. (2023), Figure 8 of Dao and Gu (2024), Figure 2 of Yang et al. (2024c), and likely others. However, our above construction shows that the sequence length doesn’t matter for memorizers with a large enough memory capacity; once the memory is large enough, the difficulty of MQAR is independent of the sequence length $T$, which may be counter to the intuitive notion of longer sequences being more difficult. Instead, models are limited by their ability to fully memorize all $P$ key-value pairs.
@@
Although our construction relied on the embedding space being large enough to allow orthonormality, trained models can still succeed without perfectly orthonormal embeddings, which we empirically show in Figure 3. We train on the MQAR task following the procedure of (Arora et al., 2023). However, we use only one test-time regression layer and only use a single size 2 convolution for constructing the keys, following our construction. We also remove the typical MLP block following the sequence layer, since our construction shows that additional nonlinearities are unnecessary for this synthetic associative recall task. We fix our key dimensions ${D_{k}}$ to be the same as the embedding size, $d_{\text{model}}=64$. We compare two test-time regression architectures, linear attention and unweighted recursive least squares (RLS) of Equation 3. We do not use gated variants of linear attention since this task does not require any forgetting; the cue-response mapping is unchanged throughout the sequence. As shown by our previous construction, even linear attention is able to solve the task perfectly when the appropriate keys are constructed. In the case where $d_{\text{model}}=P=64$, linear attention solves MQAR perfectly, independent of sequence length as predicted. Once the number of keys increases to $P=128$, linear attention can no longer perfectly solve MQAR, since it is impossible for 128 keys to be orthogonal in a $d_{\text{model}}=64$ dimensional space. In contrast, an unweighted RLS layer improves upon linear attention by accounting for the key covariance matrix $\mathbf{K}{}^{\top}\mathbf{K}$, allowing it to better test-time memorize the cue-response map, even when $P>d_{\text{model}}$. As a result, RLS is able to solve MQAR regardless of the sequence length.
@@
We have presented a unifying framework that views sequence models with associative memory as test-time regressors. Each architecture emerges naturally by specifying three aspects: the relative importance of each association, the function class, and the optimization algorithm. Through this lens, we derived linear attention variants as approximations to recursive least squares, online learners as first-order least-squares solvers, and softmax attention as a local polynomial regressor. This perspective provided principled explanations for architectural choices like QKNorm and derived higher-order generalizations of attention that go beyond pairwise interactions. We also demonstrated that for multi-query associative recall (MQAR), a single short convolution and a sequence layer suffices, in contrast to typical approaches that use multiple layers and unnecessary feedforward blocks (Arora et al., 2023).

### v3

@@ Abstract
resulting in an offset softmax attention as the output: $m_{1,\,t}(\mathbf{q})=\mathbf{M}^{[0]}=\bar{\mathbf{v}}-\mathbf{M}^{[1]}\bar{\mathbf{d}}$. This alternate form comes from the fact that if we already know $\mathbf{M}^{[1]}$, then we can solve the $p=0$ case with an alternate regression problem with preprocessed data $\{(\mathbf{k}_{i},\mathbf{v}_{i}-\mathbf{M}^{[1]}\mathbf{d}_{i})\}_{i=1}^{t}$, resulting in $\mathbf{M}^{[0]}$ being a normalized weighted sum of $\mathbf{v}_{i}-\mathbf{M}^{[1]}\mathbf{d}_{i}$. Unlike standard softmax attention (the $p=0$ case), which only considers the interaction between keys and queries, the $p=1$ case in Equation 52 also accounts for the covariance between keys through the inverse term, similar to how linear regression improves upon linear attention from before.
@@
Unfortunately, directly computing Equation 52 for $\mathbf{q}\in\{\mathbf{q}_{1},\ldots,\mathbf{q}_{T}\}$ in parallel to fully leverage hardware-acceleration is impractical for all but the smallest sequence problems. Materializing all $T$ prefix set of input tokens $\{\mathbf{X}_{t}\}_{t=1}^{T}$ requires $\mathcal{O}\left(T^{2}{D_{k}}\right)$ space, and materializing all $T$ prefix weighted sums of outer products $\{\mathbf{X}_{t}{}^{\top}\mathbf{W}_{t}\mathbf{X}_{t}\}_{t=1}^{T}$ requires $\mathcal{O}\left(T{D_{k}}^{2}\right)$ space. Computing $\{\mathbf{X}{}^{\top}_{t}\mathbf{W}_{t}\allowbreak\mathbf{X}_{t}\}_{t=1}^{T}$ requires $\mathcal{O}\left(T^{2}{D_{k}}\right)$ time, and solving a batch of $T$ linear systems requires $\mathcal{O}\left(T{D_{k}}^{3}\right)$ time. These challenges makes the current form computationally impractical for practical sequence tasks. Developing hardware-aware optimizations to compute Equation 52 similar to FlashAttention (Dao et al., 2022) is a promising future direction.
Section 5 Online regression with sequence layers
We now empirically illustrate that the sequence layers discussed so far do in fact perform regression implicitly via a single forward pass. We generate a sequence of keys according to a non-stationary, switching, autoregressive process:
@@
with ${D_{k}}=64$ and $T=256$. Each $\bm{\varepsilon}_{t}\sim\operatorname{Normal}\left(0,\bm{\Sigma}\right)$ where $\bm{\Sigma}$ has unit diagonals and non-zero off-diagonals. The choice of a non-identity $\Sigma$ creates instantaneous correlation between the dimensions of the keys so that $\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t}\neq\mathbf{I}$. We set the value vectors to be $\mathbf{v}_{t}=\mathbf{k}_{t+1}/\lVert\mathbf{k}_{t+1}\rVert_{2}$, such that the dataset $\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{T}$ imitates a next-token prediction task. We make one-step-ahead predictions by setting $\mathbf{q}_{t}=\mathbf{k}_{t+1}$ so that $\mathbf{y}_{t}={m}_{t}(\mathbf{q}_{t})={m}_{t}(\mathbf{k}_{t+1})$ is the test-time regressor’s out-of-sample prediction of the next timestep, conditioned on the prefix $\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{t}$. To solve this online regression problem, we simply compute a single forward pass through each sequence layer.
We evaluate layers derived both from parametric regression and non-parametric regression. For our parametric regression layers we choose linear attention (Equation 2) and recursive least squares (Equation 11) to show the effects of accounting for correlations between keys. For our non-parametric regression layers, we chose softmax attention (Equation 1) and our $p=1$ generalization of softmax attention based on local linear regression (Equation 52). We also show the naive one-step forecaster baseline where we simply use $\mathbf{v}_{t-1}$ as the prediction given $\mathbf{k}_{t}$.
@@
We plot the one-step out-of-sample test loss $\lVert\mathbf{v}_{t+1}-{m}_{t}(\mathbf{k}_{t+1})\rVert_{2}^{2}$ in Figure 2. During the initial $T/4$ timesteps, the dynamical system has a faster decay time with more rapid changes, making it less predictable. In the latter $43T/$ 4 timesteps, the system slows down and is more predictable. However, unweighted recursive least squares is unable to adapt to this change since it cannot discount datapoints from before the transition; as a result, it maintains a high error throughout the latter timesteps. In contrast, the nonparametric regressors are able to adapt quickly to the transition since earlier keys are dissimilar to the later keys. Interestingly, increasing our nonparametric softmax regressor from $p=0$ to $p=1$ allows the model to better adapt to both the initial fast-changing period and the sudden regime change. As expected, linear attention performs the worst since $\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t}\neq\mathbf{I}$ in our dynamical system [^4] We used a standard output normalization for linear attention (Qin et al., 2022) to prevent the output from scaling linearly with the sequence length..
@@
Section 6 Constructing effective key-value pairs for next-token recall
@@
Having shown that models derived via regression are able to perform associative recall, we now discuss the importance of constructing test-time key-value pairs that are pertinent to the task at hand. This echoes an unchanging principle in machine learning: a well-designed model is only as effective as the data it processes.
Historically, query-key-value sequences were constructed via a linear projection on the corresponding input $\mathbf{x}_{t}$ at that timestep, e.g. $\mathbf{k}_{t}=\mathbf{W}_{K}\mathbf{x}_{t}$ (Vaswani et al., 2017). Recent work on recurrent models, starting with Fu et al. (2022), identified the importance of performing a “short convolution” in addition to the per-timestep projection:
$\mathbf{k}_{t}=\sum_{i=0}^{K-1}w_{i}\mathbf{W}_{K}\mathbf{x}_{t-i}$ (56)
where $K$ is the length of the short convolution filter. The short convolution imitates the induction heads behavior learned by self-attention layers near the start of the neural network, which allow Transformers to “look back” at previous tokens. Setting $w_{0}=i$ and $w_{j}=0$ for $j\neq 0$ in Equation 56 enables other sequence layers to also “look back” at the token from timestep $i$.
The short convolution is a crucial component of purely recurrent language models; removing the short convolution results in severe performance drops (Yang et al., 2024b, c, a, Sun et al., 2024). Including a short convolution before each self-attention layer can even improve the performance of Transformers, by freeing the earlier self-attention layers from having to learn the induction head behavior (Xu et al., 2024). In this section, we provide a partial explanation for why this short convolution is so crucial, through our perspective of associative memory. Specifically, we show that this short convolution allows the sequence layer to memorize and retrieve bigram-like key-value pairs, which has been shown to be important for language modeling (Arora et al., 2023a).
One short conv is all you need for next-token recall.
@@
Consider a standard associative recall task introduced by Arora et al. (2023a), in which a model must be able to recall multiple key-value pairings, known as multi-query associative recall (MQAR). In this task, there are $P$ pairs of unique and consistent cue-response [^5] We use the “cue” and “response” terminology to avoid confusion with “keys” and “values” in the context of test-time regression. tokens $\{(\mathbf{u}_{j},\mathbf{v}_{j})\}_{j\in[P]}$ with each cue $\mathbf{u}_{j}$ mapping one-to-one to its response $\mathbf{v}_{j}$. The model is given a contextual sequence of randomly drawn, possibly repeated, cue-response pairs followed by a previously seen cue: $(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T-1},\mathbf{x}_{T},\mathbf{x}_{T+1})=(\mathbf{u}_{i_{1}},\mathbf{v}_{i_{1}},\ldots,\mathbf{u}_{i_{T/2}},\mathbf{v}_{i_{T/2}},\mathbf{u}_{i_{j}})$. $\mathbf{u}_{i_{j}}$ is a cue that appeared earlier (the $j$th pair) and the model is expected to output the paired response $\mathbf{v}_{i_{j}}$ at timestep $T+1$. This benchmark abstracts the “Hakuna Matata“ example “Hakuna Matata, it means no worries for the rest of your days. Hakuna           ” from Arora et al. (2023a), which we also discussed in Section 1.
We now show that a single test-time regression layer (e.g. as simple as linear attention) with one short convolution is sufficient to solve MQAR, without any parameters other than the embeddings, as long as it is provided with the appropriate key-value pairs to regress.
To solve this task, it suffices to test-time memorize all the bigram pairing $\{(\mathbf{u}_{j},\mathbf{v}_{j})\}$. Given $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$, we causally construct our test-time regression dataset as follows. Let the keys be constructed via a short convolution $\mathbf{k}_{t}=w_{0}\mathbf{x}_{t}+w_{1}\mathbf{x}_{t-1}=\mathbf{x}_{t-1}$ where $w_{0}=0,w_{1}=1$. Let the queries and values be the same as the inputs: $\mathbf{q}_{t}=\mathbf{v}_{t}=\mathbf{x}_{t}$. In other words, our AR sequence layer will first memorize $\{(\mathbf{k}_{t},\mathbf{v}_{t})\}_{t=1}^{T+1}=\{(\mathbf{x}_{t-1},\mathbf{x}_{t})\}_{t=1}^{T+1}$ into its associative memory map ${m}_{T+1}$. Then we retrieve from ${m}_{T+1}$ using $\mathbf{q}_{T+1}=\mathbf{x}_{T+1}=\mathbf{u}_{i_{j}}$, which has previously appeared in the sequence of cues as the $i_{j}$th bigram. The output is then
$\mathbf{y}_{T+1}={m}_{T+1}(\mathbf{q}_{T+1})={m}_{T+1}(\mathbf{u}_{i_{j}}).$ (57)
For simplicity, suppose ${m}_{T+1}$ is implemented via linear attention, following Equation 2. Then
$\mathbf{y}_{T+1}={m}_{T+1}(\mathbf{u}_{i_{j}})=\left(\sum_{t=1}^{T+1}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}\right)\mathbf{u}_{i_{j}}=\sum_{t=1}^{T+1}\mathbf{x}_{t}\mathbf{x}_{t-1}{}^{\top}\mathbf{u}_{i_{j}}.$ (58)
When the embedding space is large enough, we can set the embeddings such that all tokens are orthonormal. Then $\mathbf{x}_{t-1}{}^{\top}\mathbf{u}_{i_{j}}$ is nonzero only when $\mathbf{x}_{t-1}=\mathbf{u}_{i_{j}}$, simplifying the output to a sum of only tokens that are followed by $\mathbf{u}_{i_{j}}$. Since the cue-response map is one-to-one, the remaining tokens are all $\mathbf{v}_{i_{j}}$, producing the output $\mathbf{y}_{T+1}\propto\mathbf{v}_{i_{j}}$, solving the MQAR task.
@@
Typical evaluations of models on the MQAR task look at model performance with respect to the length of the sequence, possibly varying the model capacity. A few examples include Figure 2 of Arora et al. (2023a), Figure 8 of Dao and Gu (2024), Figure 2 of Yang et al. (2024c), and likely others. However, our above construction shows that the sequence length doesn’t matter for memorizers with a large enough memory capacity; once the memory is large enough, the difficulty of MQAR is independent of the sequence length $T$, which may be counter to the intuitive notion of longer sequences being more difficult. Instead, models are limited by their ability to fully memorize all $P$ key-value pairs.
@@
Although our construction in Equation 58 relied on the embedding space being large enough to allow orthonormality, trained models can still succeed without perfectly orthonormal embeddings, which we empirically show in Figure 3. We train on the MQAR task following the procedure of (Arora et al., 2023a). However, we use only one test-time regression layer and one length-2 convolution (for constructing the keys), following our construction. We also remove the typical MLP block following the sequence layer, since our construction shows that additional nonlinearities are unnecessary for this synthetic associative recall task. We fix our key dimensions ${D_{k}}$ to be the same as the embedding size, $d_{\text{model}}=64$. We compare two test-time regression architectures, linear attention and unweighted recursive least squares (RLS) of Equation 11. We do not use gated variants of linear attention since this task does not require any forgetting; the cue-response mapping is unchanged throughout the sequence. As shown by our previous construction, even linear attention is able to solve the task perfectly when the appropriate keys are constructed. In the case where $d_{\text{model}}=P=64$, linear attention solves MQAR perfectly, independent of sequence length as predicted. Once the number of keys increases to $P=128$, linear attention can no longer perfectly solve MQAR, since it is impossible for 128 keys to be orthogonal in a $d_{\text{model}}=64$ dimensional space. In contrast, an unweighted RLS layer (Equation 11) improves upon linear attention by accounting for the key covariance matrices $\{\mathbf{K}_{t}{}^{\top}\mathbf{K}_{t}\}$, allowing it to better test-time memorize the cue-response map, even when $P>d_{\text{model}}$. As a result, RLS is able to solve MQAR regardless of the sequence length.
Section 7 Related works
Our test-time regression framework was derived by formalizing the notion of associative recall, resulting in sequence layers that first memorize then retrieve. Designing neural networks motivated with associative memory has a long history, dating back as early as the works of Hopfield (1982), Willshaw et al. (1969),Willshaw (1989),Kohonen (1972), and Hinton and Anderson (1989). Many more contemporary works have also sought to understand and derive neural network layers from the perspective of memory, drawing inspiration from both neuroscience and dynamical systems (Bietti et al., 2023, Krotov and Hopfield, 2016, 2021, Ramsauer et al., 2021, Millidge et al., 2022, Behrouz et al., 2024). Indeed, even the feedforward layers of an MLP can be seen as a persistent associative memory (Sukhbaatar et al., 2019, Zhang et al., 2025, Bietti et al., 2023, Nichani et al., 2024, Cabannes et al., 2023).
The ability to reference and retrieve past tokens for further computation is believed to be key to in-context learning and zero-shot/few-shot learning in large language models (Olsson et al., 2022, Dong et al., 2024). There have been many attempts at understanding how sequence models are able to learn in-context (Garg et al., 2022, Zhang et al., 2023b). Some works have shown that Transformers, due to their self-attention layers, learn linear functions in-context by learning to perform gradient descent, where the iterates are updated per layer (Oswald et al., 2023, Ahn et al., 2023). This phenomenon is sometimes known as mesa-optimization (von Oswald et al., 2023) where the model learns to solve an auxiliary optimization problem in order to minimize its training objective. However, sometimes these intermediate optimization processes are actually implicitly in the architecture itself, as we showed in Equation 24 where each linear attention layer (and its gated variant) explicitly performs one step of gradient descent. Note that this behavior is different from the class of recurrent models that perform stochastic gradient descent within the layer, as defined by Equation 26.
@@
We have presented a unifying framework that derives sequence models that can explicitly perform associative recall via test-time regression. Each architecture emerges naturally by specifying three aspects: the relative importance of each association, the function class, and the optimization algorithm. Through this lens, we derived linear attention and its variants as approximations to the linear least squares solution, online learning layers and fast-weight layers as first-order least-squares solvers, and softmax attention as a local constant regressor. This perspective provided principled explanations for architectural choices like QKNorm and derived higher-order generalizations of attention that go beyond pairwise interactions. We also demonstrated that for multi-query associative recall (MQAR), a single short convolution and a sequence layer suffices, in contrast to typical approaches that use multiple layers and unnecessary feedforward blocks (Arora et al., 2023a).